# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

### v1

@@ Abstract.
Spatialized vibrotactile feedback systems deliver tactile information by placing multiple vibrotactile actuators on the body. As increasing numbers of actuators are required to adequately convey information in complicated applications, haptic designers find it difficult to create such systems due to limited scalability of existing toolkits. We propose VibraForge, an open-source vibrotactile toolkit that supports up to 128 vibrotactile actuators. Each actuator is encapsulated within a self-contained vibration unit and driven by its own microcontroller. By leveraging a chain-connection method, each unit receives independent vibration commands from a control unit, with fine-grained control over intensity and frequency. We also designed a GUI Editor to expedite the authoring of spatial vibrotactile patterns. Technical evaluations show that vibration units reliably reproduce audio waveforms with low-latency and high-bandwidth data communication. Case studies of phonemic tactile display, virtual reality fitness training, and drone teleoperation demonstrate the potential usage of VibraForge within different domains.
Haptics, Vibrotactile, Toolkits, Wearable Devices
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Haptic devicesccs: Human-centered computing User interface toolkits
Figure 1. VibraForge is a scalable, open-source vibrotactile toolkit that supports the creation of spatialized vibrotactile feedback systems. Leveraging a chain-connection method, one control unit in the system can independently control up to 128 vibration units. Applications of the toolkit are shown on the right.
@@
An overview of the VibraForge toolkit, that has control and vibration units, uses a chain-connection topology, and supports the creation of multi-actuator devices.
@@ 1. Introduction
Tactile feedback conveys sensory information to the human body through the sense of touch (El Saddik et al., 2011). Vibrotactile feedback, which produces tactile sensations via actuators vibrating on the skin, has gained much popularity because it is compact, affordable, and accessible (Choi and Kuchenbecker, 2013). It has been widely adopted in various domains, such as virtual reality (VR) (Wang and et al., 2021; Kaul and Rohs, 2017; bHaptics, 2024), robotics (Pacchierotti and Prattichizzo, 2024; Xia et al., 2022; Aggravi et al., 2018), rehabilitation (Bark et al., 2014; Prabhu et al., 2020; Lindeman et al., 2006), and accessibility (Witteveen et al., 2015; Jung et al., 2022; Reed et al., 2018).
@@
To provide adequate feedback to users during complex tasks, spatialized vibrotactile feedback systems have been proposed (Lindeman et al., 2004; Louison et al., 2017). These systems utilize actuator positions as an additional channel to enhance information transmission. For example, Wang et al. designed a facial mask with 36 eccentric rotating mass (ERM) actuators to study the perception of localization and orientation of objects in VR  (Wang and et al., 2021). The bHaptics TactSuit X40 (bHaptics, 2024) utilized 40 ERMs on a vest to create localized feedback for immersive VR gaming experiences. The bHaptics TactSuit was also used by Xia et al. to convey hydrodynamic flow intensity during submarine teleoperation (Xia et al., 2022). Jung et al. designed a wireless haptic interface with 36 ERMs to convey detailed fingertip touch sensations on the upper arm for amputees (Jung et al., 2022). Sanchez et al. created OpenVNAVI, a navigation aid system with 128 ERMs to guide visually impaired people (Sánchez and Bohne, 2015). A 64-actuator tactile jacket was also created by Lemmens et al. to enrich movie viewing experiences (Lemmens et al., 2009). Similarly, West et al. created the Body:Suit:Score system with 60 ERMs to enhance musical experiences (West et al., 2019).
@@
Despite the great potential of spatialized vibrotactile feedback systems, previous implementations of such systems revealed scalability and expressivity limitations. Systems that depend on direct connections between actuators and microcontroller units (MCUs) exhibit limited scalability, as the number of actuators is intrinsically restricted by the available GPIOs or PWM channels on a MCU (Wang and et al., 2021; Kaul and Rohs, 2017; Reed et al., 2018). Expanding the number of actuators requires multiple MCUs, resulting in exponential increases in system complexity and communication overhead. For systems that use custom drivers and data bus protocols to drive actuators indirectly, the multi-layer connections support more actuators but compromise the vibration expressivity (LI et al., 2013; Lemmens et al., 2009; West et al., 2019). These systems cannot render complex waveforms such as audio signals due to bandwidth limits, thereby limiting the types of actuators that can be used to simple actuators like ERMs. Moreover, while various toolkits have been developed to assist vibrotactile design, they typically employ direct or multi-layer connections and face the same limitations (e.g., TECHTILE (Minamizawa et al., 2012), VITAKI (Martínez et al., 2014), and VHP (Dementyev et al., 2021)).
@@
To overcome these limitations, this work proposes VibraForge, an open-source vibrotactile toolkit that leverages a scalable, modular design to enable the rapid prototyping of vibrotactile systems. VibraForge consists of control units and vibration units. Control units are responsible for receiving vibration control commands from a Bluetooth server and sending them to vibration units. Vibration units are self-contained modules with an actuator driven by their own microcontroller. Vibration units are connected to a control unit via a chain-connection method, where a single GPIO pin on a control unit can control up to sixteen vibration units via a custom UART protocol. To better support spatial pattern authoring, we also designed an accompanying GUI Editor that enables haptic designers to intuitively author multi-actuator patterns.
@@
Technical evaluations of VibraForge showed that the toolkit has a high bandwidth (200 Hz), low latency (16 ms), and can reliably render complex audio signals in real time. Through three case studies, we demonstrate that VibraForge not only supports the efficient replication of existing research prototypes such as phonemic displays (Reed et al., 2018), but also improves user performance in various applications such as virtual reality fitness gaming and robot teleoperation.
The contributions of this work are:
@@
  • an open-source hardware and software toolkit, VibraForge, that supports the design of spatialized vibrotactile feedback systems,
@@
  • technical evaluations of VibraForge, including power consumption, bandwidth, latency and acceleration, etc., and
  • case studies demonstrating replicated research prototypes and novel applications enabled by VibraForge.
@@ 2.1. Existing Vibrotactile Toolkits
Vibrotactile actuators such as eccentric rotation mass vibration motors (ERM), linear resonant actuators (LRA), voice coil actuators (VCA), or piezoelectric actuators (PIEZO) create mechanical vibrations via electromagnetic motions. Each actuator type and model has distinct voltage ratings, frequency responses, and vibration characteristics, making it challenging for designers to explore different configurations for wearable tactile devices. To assist the design process, various vibrotactile toolkits have been developed. These toolkits abstract low-level technical details and offer intuitive design interfaces, so researchers and designers can concentrate on creating vibration effects without being overwhelmed by the complexities of electronic, mechanical, and software design.
@@
While existing toolkits provide rich features to support vibrotactile design, they are limited by their scalability, particularly in the number of actuators they can support. This is because the actuators are directly controlled via GPIO pins or PWM channels on MCUs, with most offering a maximum of 16 channels. As vibrotactile feedback continues to be used for more complicated tasks and applications, these toolkits cannot meet the demands of emerging systems. Therefore, when designing VibraForge, we drew insights from previous spatialized vibrotactile feedback systems and propose a new connection method to overcome these scalability limitations.
@@ 2.2. Spatialized Vibrotactile System Design
Two connection methods have been often deployed When building spatialized vibrotactile systems (Figure 2). Some have tried maintaining a ”direct connection” to each actuator by adding more controllers to a system and thus increasing the number of PWM channels. For example, Wang et al. used PWM outputs from four Arduino Mega2560 boards to drive their facial display (Wang and et al., 2021). Similarly, Kaul et al. used four Arduino Nanos to drive 17 actuators mounted on their HapticHead display (Kaul and Rohs, 2017). Reed et al. used three custom-built audio amplifier boards to send individual waveforms to 24 actuators on a sleeve to deliver speech phonemic cues to deaf users (Reed et al., 2018). Although complex waveforms could be sent to the actuators via direct connections, communication overhead and system complexity increased significantly when the system scaled up.
[…]
@@
There is thus a trade-off between system scalability and actuator expressivity: a system that supports fine-grained control of individual actuators is difficult to scale, and vice versa. In this work, we propose a chain-connection method (Figure 2), where actuators are encapsulated into self-contained vibration units and driven by their own PCB drivers. Chains of vibration units are connected to a central MCU. Each PCB driver receives commands from the central MCU and generates PWM signals for its own actuator, thus supporting scalability. The high bandwidth and low latency control of vibration characteristics enables actuators to render complex waveforms, thus enhancing expressivity
@@ 2.3. Software for Vibration Pattern Design
Vibration pattern design is crucial in spatialized vibrotactile feedback systems. For a comprehensive review of previous design software and their public availability, please see (Terenti and Vatavu, 2023). Here we discuss the common interfaces used to design single-actuator waveforms and multi-actuator patterns.
For single-actuator design, one common method is to use predefined waveforms. Designers start with standard waveforms such as sine, square, or triangle, multiply them to generate new waveforms, and modify them using envelopes (Pezent et al., 2020). While this method provides authoring flexibility, it requires that designers have a deep understanding of how waveforms can be composed through combinations of standard signals. This makes it difficult for novice designers to use. Others have created editors that enable the direct manipulation of waveforms. With these editors, designers have access to two parallel windows to edit vibration amplitude and frequency. For example, in the Macaron Editor, designers could directly move keyframe dots to modify a waveform, making the design process more intuitive (Schneider and MacLean, 2016). This method was later adopted by several industrial software tools, including Meta Haptics Studio (Meta Platforms, 2023a).
@@
VibraForge’s GUI Editor combines single-actuator waveform design with a multi-actuator canvas to support the efficient design for spatialized systems. It provides a library of standard waveforms and envelops for composing complex waveforms. It is also compatible with existing tools such as Syntacts (Pezent et al., 2020) and VIREO (Terenti and Vatavu, 2023) to support the reuse of workflows and prior knowledge. The canvas shows a virtual layout of actuators that replicates the physical chain connections so designers can use it as reference when designing spatial patterns.
@@ 3. VibraForge Design
As the discussion of prior work highlighted, there is a trade-off between actuator expressivity and system scalability. We aimed to mitigate this trade-off by using a novel chain-connection method with custom designed vibration units and control units. This section provides an overview of the VibraForge toolkit [^1] The project resource can be found at: https://anonymous.4open.science/r/VibraForge, including its design requirements, vibration units, control units, chain-connection method, signal segmentation algorithm, and vibration pattern GUI Editor.
@@ 3.1. Design Requirements
VibraForge’s design requirements were drawn from design challenges discussed in prior work on wearable spatial vibrotactile design or were identified by the authors based on prior prototyping experiences.
From spatial vibrotactile design, three design requirements were proposed:
  • DR1: Scalability. The toolkit should support large quantities of actuators. Designers should be able to use different types of actuators such as ERM, LRA, and VCA.
@@
  • DR3: Usability. The toolkit should provide an intuitive design interface for authoring vibrotactile spatial patterns. To leverage designers’ prior design knowledge, the toolkit should be compatible with existing vibrotactile waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a).
From wearable design, two design requirements were proposed:
  • DR4: Flexibility. The toolkit should enable designers to modify settings, reprogram behaviors, and physically rearrange components without difficulty. This requirement was derived from design challenges faced by West et al. when attempting to relocate or replace components that were sewn onto a garment (West et al., 2019).
  • DR5: Portability. Devices built with the toolkit should be portable and powered by batteries, communicate with external devices via wireless protocols, and use lightweight components to ensure a comfortable wearing experience.
@@ 3.2. Vibration Units
The vibrations units were designed as self-contained modules that could drive actuators by themselves. Each vibration unit consisted of a vibrotactile actuator, a custom designed PCB, and 3D printed components (Figure 3). We designed three versions of the vibration unit to support a range of actuators (DR1).
Figure 3. (Left) Exploded views of the small vibration unit for X-axis LRAs, including the cap, the PCB, the PCB mount, the enclosure, and the bottom ring. Z-axis LRAs shared most components except the enclosure. (Right) Exploded view of the large vibration unit for VCAs with a similar structure. Note that the bottom ring was placed underneath the garment to attach the unit.

### v2

@@ Abstract.
Spatialized vibrotactile feedback systems deliver tactile information by placing multiple vibrotactile actuators on the body. As increasing numbers of actuators are required to adequately convey information in complicated applications, haptic designers find it difficult to create such systems due to limited scalability of existing toolkits. We propose VibraForge, an open-source vibrotactile toolkit that supports up to 128 vibrotactile actuators. Each actuator is encapsulated within a self-contained vibration unit and driven by its own microcontroller. By leveraging a chain-connection method, each unit receives independent vibration commands from a control unit, with fine-grained control over intensity and frequency. We also designed a GUI Editor to expedite the authoring of spatial vibrotactile patterns. Technical evaluation showed that vibration units reliably reproduced audio waveforms with low-latency and high-bandwidth data communication. Case studies of a phonemic tactile display, virtual reality fitness training, and drone teleoperation demonstrated the potential usage of VibraForge within different domains. A usability study with non-expert users highlighted the low technical barrier and customizability of the toolkit.
Haptics, Vibrotactile Feedback, Haptic Toolkit, Multi-Actuator Devices, Wearable Devices
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714273isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Haptic devicesccs: Human-centered computing User interface toolkits
Figure 1. VibraForge is a scalable, open-source vibrotactile toolkit that supports the creation of spatialized vibrotactile feedback systems. Leveraging a chain-connection method, one control unit in the system can independently control up to 128 vibration units. Applications of the toolkit include phonemic tactile displays, VR fitness training, and drone teleoperation.
@@
An overview of the VibraForge toolkit, which has control units and vibration units, uses a chain-connection topology, and supports the creation of spatialized vibrotactile feedback systems.
@@ 1. Introduction
Tactile feedback conveys sensory information to the human body through the sense of touch (Weber, 1996). Vibrotactile feedback, which produces tactile sensations via actuators vibrating on the skin, has gained much popularity because it is compact, affordable, and accessible, as summarized by Choi et al. (Choi and Kuchenbecker, 2013). It has been widely adopted in various domains, such as virtual reality (VR) (Wang et al., 2021; Kaul and Rohs, 2017; bHaptics, 2024), robotics (Pacchierotti and Prattichizzo, 2024; SNAME, 2022; Aggravi et al., 2018), rehabilitation (Bark et al., 2014; Prabhu et al., 2020; Lindeman et al., 2006), and accessibility (Witteveen et al., 2015; Jung et al., 2022; Reed et al., 2018).
@@
To provide adequate feedback to users during complex tasks, spatialized vibrotactile feedback systems have been proposed, such as a battle simulation suit by Lindeman et al. (Lindeman et al., 2004) and directed movement sleeves by Louison et al. (Louison et al., 2017). These systems utilize actuator positions as an additional channel to enhance information transmission. For example, Wang et al. designed a facial mask with 36 eccentric rotating mass (ERM) actuators to study the perception of localization and orientation of objects in VR  (Wang et al., 2021). The bHaptics TactSuit X40 (bHaptics, 2024) utilized 40 ERMs on a vest to create localized feedback for immersive VR gaming experiences. The bHaptics TactSuit was also used by Xia et al. to convey hydrodynamic flow intensity during submarine teleoperation (SNAME, 2022). Jung et al. designed a wireless haptic interface with 36 ERMs to convey detailed fingertip touch sensations on the upper arm for amputees (Jung et al., 2022). Sanchez et al. created OpenVNAVI, a navigation aid system with 128 ERMs to guide visually impaired people (Sánchez and Bohne, 2015). A 64-actuator tactile jacket was also created by Lemmens et al. to enrich movie viewing experiences (Lemmens et al., 2009). Similarly, West et al. created the Body:Suit:Score system with 60 ERMs to enhance musical experiences (West et al., 2019).
@@
Despite the great potential of spatialized vibrotactile feedback systems, previous implementations of such systems revealed scalability and expressivity limitations. Systems that depend on direct connections between actuators and microcontroller units (MCUs) exhibit limited scalability, as the number of actuators is intrinsically restricted by the available GPIOs or PWM channels on a MCU (Wang et al., 2021; Kaul and Rohs, 2017; Reed et al., 2018). Expanding the number of actuators requires multiple MCUs, resulting in exponential increases in system complexity and communication overhead. For systems that use custom drivers and data bus protocols to indirectly drive actuators, multi-layer connections support more actuators but compromise vibration expressivity (LI et al., 2013; Lemmens et al., 2009; West et al., 2019). These systems cannot render complex waveforms such as audio signals due to bandwidth limits, thereby limiting the types of actuators that can be used to simple actuators like ERMs. Moreover, while various toolkits have been developed to assist vibrotactile design, they typically employ direct or multi-layer connections and face the same limitations (e.g., TECHTILE (Minamizawa et al., 2012), VITAKI (Martínez et al., 2014), and VHP (Dementyev et al., 2021)).
@@
To overcome these limitations, this work proposes VibraForge, an open-source vibrotactile toolkit [^1] The project resource can be found at: https://github.com/huangbj16/VibraForge that leverages a scalable, modular design to enable the rapid prototyping of vibrotactile systems. VibraForge consists of control units and vibration units. Control units receive vibration control commands from a Bluetooth server and sending them to vibration units. Vibration units are self-contained modules with an actuator driven by their own microcontroller. They are connected to control units via a chain-connection method, where a single GPIO pin on a control unit can control up to sixteen vibration units via a custom UART protocol. To better support spatial pattern authoring, we also designed an accompanying GUI Editor that enables haptic designers to intuitively author multi-actuator tactile patterns.
@@
Technical evaluation of VibraForge showed that the toolkit has a high bandwidth (200 Hz), low latency (16 ms), and can reliably render complex audio signals in real time. Through three case studies, we demonstrate that VibraForge not only supports the efficient replication of existing research prototypes such as phonemic displays (Reed et al., 2018), but also improves user performance in various applications such as virtual reality fitness gaming and robot teleoperation. In a usability study, ten participants with diverse technical backgrounds learned to use the toolkit in 15 minutes and completed design tasks and a free exploration activity. The findings from this study reaffirmed VibraForge’s low technical barrier and high potential for customizability.
The contributions of this research are:
@@
  • an open-source hardware toolkit that supports the design of spatialized vibrotactile feedback systems,
  • a novel chain-connection method that addresses the gap between system scalability and feedback expressivity,
  • a software GUI editor that enables designers to intuitively author spatial vibration patterns,
@@
  • technical evaluations of VibraForge, including latency, bandwidth, vibration characteristics, etc., and
  • case studies and a usability study demonstrating the usefulness and usability of the toolkit.
[…]
@@ 2.1. Existing Vibrotactile Toolkits
Vibrotactile actuators such as eccentric rotation mass vibration motors (ERM), linear resonant actuators (LRA), voice coil actuators (VCA), or piezoelectric actuators (PIEZO) create mechanical vibrations via electromagnetic motions. Each actuator type and model has distinct voltage ratings, frequency responses, and vibration characteristics, making it challenging for designers to explore different configurations when creating wearable tactile devices. To assist their design process, various vibrotactile toolkits have been proposed. These toolkits abstract low-level technical details and offer intuitive design interfaces, so designers can concentrate on creating vibration effects without being overwhelmed by the complexities of electronic, mechanical, and software design.
@@
While existing toolkits provide rich features to support vibrotactile design, they are limited by their scalability, particularly in the number of actuators they can support. This is because the actuators are directly controlled via GPIO pins or PWM channels on MCUs, with most offering a maximum of 16 channels. As vibrotactile feedback continues to be used for more complicated tasks and applications, these toolkits cannot meet the demands of emerging systems. Therefore, when designing VibraForge, we drew insights from previous spatialized vibrotactile feedback systems and created a new connection method to overcome these scalability limitations.
@@ 2.2. Spatialized Vibrotactile System Design
Two connection methods are often used when building spatialized vibrotactile systems (Figure 2). Some have tried maintaining a ”direct connection” to each actuator by adding more microcontrollers to a system, thus increasing the number of PWM channels. For example, Wang et al. used PWM outputs from four Arduino Mega2560 boards to drive their facial display (Wang et al., 2021). Similarly, Kaul et al. used four Arduino Nanos to drive 17 actuators mounted on their HapticHead display (Kaul and Rohs, 2017). Reed et al. used three custom-built audio amplifier boards to send individual waveforms to 24 actuators on a sleeve to deliver speech phonemic cues to deaf users (Reed et al., 2018). Although complex waveforms could be sent to the actuators via direct connections, communication overhead and system complexity increased significantly when the system was scaled up.
@@
There is thus a trade-off between system scalability and actuator expressivity: a system that supports fine-grained control of individual actuators is difficult to scale, and vice versa. In this work, we propose a chain-connection method (Figure 2), where actuators are encapsulated into self-contained vibration units and driven by their own PCB drivers. Chains of vibration units are connected to a central MCU. Each PCB driver receives commands from the central MCU and generates PWM signals for its own actuator, thus supporting scalability. The high bandwidth and low latency control of vibration characteristics enables actuators to render complex waveforms, thus enhancing expressivity.
@@ 2.3. Software for Vibration Pattern Design
Vibration pattern design is crucial in spatialized vibrotactile feedback systems. For a comprehensive review of previous design software and their public availability, please see (Terenti and Vatavu, 2023). Here, we discuss the common interfaces used to design single-actuator waveforms and multi-actuator patterns.
For single-actuator design, designers often start with standard waveforms (e.g., sine, square, etc), multiply them to generate new waveforms, and modify them using envelopes, as illustrated in Syntacts (Pezent et al., 2020). While this method provides authoring flexibility, it requires that designers have a deep understanding of how waveforms can be composed through combinations of standard signals. This makes it difficult for novice designers to use. Others have created editors that enable the direct manipulation of waveforms. With these editors, designers have access to two parallel windows to edit vibration amplitude and frequency. For example, in the Macaron Editor, designers could directly move keyframe dots to modify a waveform, making the design process more intuitive (Schneider and MacLean, 2016). This method was later adopted by several industrial software tools, including Meta Haptics Studio (Meta Platforms, 2023a).
@@
VibraForge’s GUI Editor combines single-actuator waveform design with a multi-actuator canvas to support the efficient design for spatialized systems. It provides a library of standard waveforms and envelopes for composing complex waveforms. It is also compatible with existing tools such as Syntacts (Pezent et al., 2020) and VIREO (Terenti and Vatavu, 2023) to support the reuse of workflows and prior knowledge. The canvas shows a virtual layout of actuators that replicates the physical chain connections so designers can reference it while designing spatial patterns.
@@ 3. VibraForge Design
As highlighted in prior work, a trade-off exists between actuator expressivity and system scalability. To address this, we developed a novel chain-connection method with custom vibration units and control units. This section provides an overview of the VibraForge toolkit, covering its design requirements, components, chain-connection method, signal segmentation algorithm, and GUI Editor.
@@ 3.1. Design Requirements
VibraForge’s design requirements were drawn from design challenges discussed in prior work on spatial vibrotactile and wearable design, or identified by the authors based on prior prototyping experiences.
  • DR1: Scalability. The toolkit should support large quantities of actuators such as ERM, LRA, and VCA.
@@
  • DR3: Usability. The toolkit should provide an intuitive design interface for authoring vibrotactile spatial patterns that leverages designers’ prior design knowledge and is compatible with existing vibrotactile waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a).
  • DR4: Flexibility. The toolkit should enable designers to modify settings, reprogram behaviors, and physically rearrange components without difficulty, derived from design challenges faced by West et al. (West et al., 2019).
  • DR5: Portability. Devices built with the toolkit should be powered by batteries, communicate with external devices via wireless protocols, and use lightweight components to ensure a comfortable wearing experience.
@@ 3.2. Vibration Units
The vibrations units were self-contained modules comprised of a vibrotactile actuator, a custom designed PCB, and 3D printed components (Figure 3). Three versions were designed to support various actuators (DR1).
Figure 3. (Left) Exploded views of the small vibration unit for X-axis LRAs, including the cap, the PCB, the PCB mount, the enclosure, and the bottom ring. Z-axis LRAs shared most components except for the enclosure. (Right) Exploded view of the large vibration unit for VCAs with a similar structure. Note that the bottom ring was placed underneath the garment to attach the unit.